Problem 4.2: An outbreak of African swine fever, a highly contagious disease thats been called pig Ebola, is ravaging Asia
Answer #1

Salution: Given that p: - PM-P) Plt) - viguses at time t where aro and mxo are constant pig will will die when Plt) a of 5 -s m² sec? odo at (tana)? (m 2 » r m setodo = 2 sdo at (99)? segon 7+70 = out a 7+ you =0(= we have - P- m 2 = m/ tano - tano- P(-12= 7 [cittan (mat+c] PCO) = (1+tano ) - Po →ittance Pro > tarc = 220- > c= tanil 20-1) Pig dies when Plt) 00 and this hwhile Pc = 1000, M: 100, a01001, Value of -103 TP 2 31-0) = 20 (c) Joo 10 catan ( 2 P 12- tan(2x loge is 100 = tant! (20-12